### v2

@@ 1 Introduction
Single object tracking (SOT) is a long-standing topic in computer vision. Over the past years, tracking has mainly relied on 2D image data, witnessing many advanced techniques, such as Siamese region proposal network [16, 3, 35] and one-stream framework [5, 39]. With the rapid development of LiDAR sensors recently, increasing efforts [10, 25, 42] have been devoted to 3D SOT on point clouds. Current methods [25, 41, 44, 36, 37] mostly follow appearance matching paradigm that utilizes template target information to enhance target-specific feature of search region for subsequent target localization. Despite demonstrated success, incomplete and textureless point clouds introduce insufficient appearance information, limiting the application of appearance matching paradigm.
@@
M2Track [42] introduces a motion paradigm specialized for 3D SOT, achieving significant improvements in both performance and efficiency over appearance matching baseline P2B [25], as shown in Fig. 1. It first segments foreground points corresponding to tracked target from the cropped point clouds of two consecutive frames, and then infers 4-DOF [^1] The length, width and height of target bounding box are given in the first frame, and are unchanged across all frames in a point cloud sequence. relative motion of the target between the two frames. The relative motion is applied to the target position in previous frame to locate the target in current frame. Through a comprehensive analysis, we find that the effective utilization of background points can assist in target localization, which has also been explored and validated in CXTrack [36]. Therefore, a question is naturally raised: Is it possible to infer target’s 4-DOF relative motion directly from cropped point clouds of two consecutive frames?
@@
Motivated by these exploratory experiments, we casts tracking task as a direct inference of target relative motion and propose to learn detailed information changes between consecutive point clouds in a part-to-part manner to enhance motion cues. We thereby introduce a novel tracking framework, termed P2P that effectively models target-related motion cues in each part of two-frame point cloud flow. Considering the disordered nature of LiDAR point clouds, a direct part-to-part fusion between the consecutive point clouds at the input level seems to be infeasible. Towards this problem, we leverage point and voxel representations to form ordered point cloud features with implicit and explicit spatial structure information, respectively. Therefore, we propose P2P-point and P2P-voxel, which fuse each corresponding part between the spatial structure of consecutive point clouds in feature-level. Benefiting from part-to-part fusion, fine-grained motion cues can be generated, thereby guiding accurate tracking for point cloud objects. Compared with M2Track [42], both of P2P-point and P2P-voxel get rids of auxiliary modules such as segmentation and two-stage motion refinement, while achieving better performance and higher speed. Comprehensive experiments are conducted on KITTI [9], NuScenes [2] and Waymo Open Dataset (WOD) [27]. The results confirm that our proposed P2P is a strong tracking framework that achieves leading performance while maintaining high efficiency.
@@
  • We introduce P2P-point and P2P-voxel models, which explpre fine-grained motion cues by incorporating implicit and explicit part-to-part motion modeling via point- and voxel-based representation, respectively.
@@ 2 Related Work
In the context of 2D SOT, it often makes sense to decompose the tracking problem into classification and regression sub-tasks. The former focuses on localizing the target center, while the latter is dedicated to estimate the target size. Early correlation filter based tracking methods, such as KCF [12] and ECO [6], face challenges in performing both of the two sub-tasks simultaneously. Siamese network based appearance matching methods have proven effective in addressing these challenges, thus establishing a solid foundation for advancements in this field. Prominent trackers include SiamFC [1], SiamRPN [7], SiamCAR [4], STARK [38], MixFormer [5], and other notable variants [40, 3, 30, 34, 8]. Recently, appearance matching techniques have inspired many excellent works on 3D single object tracking on point clouds.
@@
The pioneering work within the context of 3D SOT is SC3D [10]. It generates a set of candidate 3D bounding boxes using a Kalman filter, and selects one box with the highest similarity to the given template target as predicted result. However, SC3D is not an end-to-end framework and fails to run in real-time due to the exhaustive candidates. Next work, P2B [25] presents a 3D region proposal network leveraging VoteNet [24]. It significantly enhances tracking performance while enabling a real-time speed. Based on this strong baseline, many follow-ups [26, 32, 41, 44, 20, 21, 36, 37, 19] have emerged. Taking recent methods as examples, CXTrack [36] designs a target-centric transformer network to explore contextual information, and MBPTrack [37] improve CXTrack with a memory network and a box-prior localization network. Following recently popular one-stream tracking framework [5, 39], SyncTrack [19] synchronizes feature extraction and matching.
@@ 3.1 Overview
In this section, we present our tracking framework P2P, along with two tracking versions, i.e., P2P-point and P2P-voxel. To provide a better understanding, we first introduce the definition of 3D single object tracking task in Sec. 3.2, followed by a thorough revisit of the existing motion tracking paradigm in Sec. 3.3. Then, we detail the proposed P2P framework and the corresponding tracking models using both point- and voxel-based representations in Sec. 3.43.5 and 3.6, respectively.
@@ 3.3 Revisit Motion Paradigm
$[\Delta x_{t},\Delta y_{t},\Delta z_{t},\Delta\theta_{t}]=F_{r}(F_{m}(F_{s}(\mathcal{P}_{t-1}^{crop},\mathcal{P}_{t}^{crop}))),$ (1)
@@ 3.4 Proposed P2P Framework
$[\Delta x_{t},\Delta y_{t},\Delta z_{t},\Delta\theta_{t}]=F_{m}^{pp}(\mathcal{P}_{t-1}^{crop},\mathcal{P}_{t}^{crop}),$ (2)
where $F_{m}^{pp}$ can mathematically express our P2P framework. Its overall architecture is illustrated in Fig. 4. The framework consists of three components: feature extractor, part-to-part motion modeling, and motion prediction head. The feature extractor distinguishes and extracts spatial structure features of $\mathcal{P}_{t-1}^{crop}$ and $\mathcal{P}_{t}^{crop}$ by a weight-shared network using dual branches. The part-to-part motion modeling module first fuse each corresponding part between the spatial structures of $\mathcal{P}_{t-1}^{crop}$ and $\mathcal{P}_{t}^{crop}$, and then models target-related motion cues in each part. Finally, the prediction head infers the 4-DOF relative motion.
@@ 3.6 Version II: P2P-voxel
$\begin{array}[]{c}1\times 4\end{array}$ $\begin{bmatrix}\!1,512,1\!\\ \!1,256,1\!\\ \!1,128,1\!\\ \!1,4,1\!\\ \end{bmatrix}\times 1$ $\begin{array}[]{c}1\times 4\end{array}$ $\begin{bmatrix}\!1\times 1,512,1\!\\ \!1\times 1,256,1\!\\ \!1\times 1,128,1\!\\ \!1\times 1,4,1\!\\ \end{bmatrix}\times 1$
@@ 3.7 Effectiveness Analysis
As a result, more fine-grained motion cues in each part are modeled, guiding more accurate and robust motion prediction. Benefiting from this, our method works in an one-stage manner without extra segmentation and two-stage refinement modules used in M2Track [42], while demonstrating state-of-the-art performance and high speed.
P2P-point v.s. P2P-voxel. P2P-voxel has certain advantages in fuse each corresponding part between the spatial structures of consecutive point clouds, compared to P2P-point. Through voxel representation, P2P-voxel effectively preserves the spatial structure of point clouds, enabling explicit part-to-part fusion. In contrast, P2P-point relies on point representation, embedding the spatial structure of point clouds into semantic features through pooling operations, which allows for an implicit part-to-part fusion between point cloud spatial structures. Refer to Fig. 6 for details.
@@ 4 Experiment
4.1 Experiment Setting
Model Input. We define search regions for two consecutive frames at timestamps $t-1$ and $t$ by centering on the previous target bounding box within a 3D space. The ranges are defined as [(-4.8,4.8),(-4.8,4.8),(-1.5,1.5)] and [(-1.92,1.92),(-1.92,1.92),(-1.5,1.5)] to contain relevant points for cars and humans, respectively. To be a fair comparison, we follow existing methods [25, 42] to sub-sample 1024 points using farthest point sampling [23] within the derived search regions as inputs for P2P-point. For P2P-voxel, we do not down-sample point clouds. Instead, we divide the search regions into 128$\times$128$\times$20 voxel grids, which serve as inputs for the model. More model details can be found in Tab. 1 and supplementary material.
@@ 4.1 Experiment Setting
Datasets. We conduct comprehensive experiments on three widely used datasets, including KITTI [9], NuScenes [2] and Waymo Open Dataset (WOD) [27]. KITTI consists of 21 training and 29 test point cloud sequences, respectively. Due to the unavailability of test labels, the training sequences are divided into training and testing sets. Compared to KITTI, NuScenes and WOD provide more challenging and large-scale scenes. NuScenes contains 700 and 150 scenes for training and testing, while WOD comprises 1121 tracklets categorized into easy, medium and hard sub-sets based on the point cloud sparsity.

### v3

@@ 1 Introduction
Single object tracking (SOT) is a long-standing topic in computer vision. Over the past years, tracking has mainly relied on 2D image data, witnessing many advanced techniques, such as Siamese region proposal network [18, 23, 3, 42, 47] and one-stream framework [6, 51]. With the rapid development of LiDAR sensors recently, increasing efforts [11, 31, 54] have been devoted to 3D SOT on point clouds. Current methods [31, 53, 4, 33, 27, 14] mostly follow appearance matching paradigm that utilizes template target information to enhance target-specific feature of search region for subsequent target localization. Despite demonstrated success, incomplete and textureless point clouds introduce insufficient appearance information, limiting the application of appearance matching paradigm.
@@
M2Track [54] introduces a motion paradigm specialized for 3D SOT, achieving significant improvements in both performance and efficiency over appearance matching baseline P2B [31], as shown in Fig. 1. It first segments foreground points corresponding to tracked target from the cropped point clouds of two consecutive frames, and then infers 4-DOF [^1] The length, width and height of target bounding box are given in the first frame, and are unchanged across all frames in a point cloud sequence. relative motion of the target between the two frames. The relative motion is applied to the target position in previous frame to locate the target in current frame. Through a comprehensive analysis, we find that the effective utilization of background points can assist in target localization, which has also been explored and validated in CXTrack [48]. In addition, M-Vanilla [55] further demonstrates the potential of the motion paradigm by removing additional auxiliary modules. Therefore, a question is naturally raised: Is it possible to infer target’s 4-DOF relative motion directly from cropped point clouds of two consecutive frames?
@@
Motivated by these exploratory experiments, we cast tracking task as a direct inference of target relative motion. To form accurate and robust motion cues, we propose to learn detailed information changes between each corresponding portion of consecutive point clouds, which we define as part-to-part motion modeling. We thereby introduce a novel tracking framework, termed P2P that effectively models target-related motion cues in each part of two-frame point cloud flow. Considering the disordered nature of LiDAR point clouds, a direct part-to-part fusion between the consecutive point clouds at the input level seems to be infeasible. Towards this problem, we leverage point and voxel representations to form ordered point cloud features with implicit and explicit spatial structure information, respectively. Therefore, we propose P2P-point and P2P-voxel, which fuse each corresponding part between the spatial structure of consecutive point clouds in feature-level. Benefiting from part-to-part fusion, fine-grained motion cues can be generated, thereby guiding accurate tracking for point cloud objects. Compared with M2Track [54], both of P2P-point and P2P-voxel get rids of auxiliary modules such as segmentation and two-stage motion refinement, while achieving better performance and higher speed. Comprehensive experiments are conducted on KITTI [10], NuScenes [2] and Waymo Open Dataset (WOD) [34]. The results confirm that our proposed P2P is a strong tracking framework that achieves leading performance while maintaining high efficiency.
@@
  • We introduce P2P-point and P2P-voxel models, which explore fine-grained motion cues by incorporating implicit and explicit part-to-part motion modeling via point- and voxel-based representation, respectively.
@@ 2 Related Work
In the context of 2D SOT, it often makes sense to decompose the tracking problem into classification and regression sub-tasks [58]. The former focuses on localizing the target center, while the latter is dedicated to estimating the target size. Early correlation filter based tracking methods, such as KCF [13] and ECO [7], face challenges in performing both of the two sub-tasks simultaneously. Siamese network based appearance matching methods have proven effective in addressing these challenges, thus establishing a solid foundation for advancements in this field. Prominent trackers include SiamFC [1], SiamRPN [8], SiamCAR [5], STARK [50], MixFormer [6], and other notable variants [52, 45, 37, 44, 9, 46, 43, 22]. Recently, appearance matching techniques have inspired many excellent works on 3D single object tracking on point clouds.
@@
The pioneering work within the context of 3D SOT is SC3D [11]. It generates a set of candidate 3D bounding boxes using a Kalman filter, and selects one box with the highest similarity to the given template target as predicted result. However, SC3D is not an end-to-end framework and fails to run in real-time due to the exhaustive candidates. Next work, P2B [31] presents a 3D region proposal network leveraging VoteNet [30]. It significantly enhances tracking performance while enabling a real-time speed. Based on this strong baseline, many follow-ups [32, 39, 53, 57, 24, 25, 48, 49, 21, 40, 26] have emerged. Taking recent methods as examples, CXTrack [48] designs a target-centric transformer network to explore contextual information, and MBPTrack [49] improve CXTrack with a memory network and a box-prior localization network. Following recently popular one-stream tracking framework [6, 51], SyncTrack [21] synchronizes feature extraction and matching.
@@ 3.1 Overview
In this section, we present our tracking framework P2P, along with two tracking versions, i.e., P2P-point and P2P-voxel. To provide a better understanding, we first introduce the definition of 3D single object tracking task in Sec. 3.2, followed by a thorough revisit of the existing motion tracking paradigm in Sec. 3.3. Then, we detail the proposed P2P framework and the corresponding tracking models using both point- and voxel-based representations in Secs. 3.43.5 and 3.6, respectively.
@@ 3.3 Revisit Motion Paradigm
$[\Delta x_{t},\Delta y_{t},\Delta z_{t},\Delta\theta_{t}]=F_{r}(F_{m}(F_{s}(\mathcal{P}_{t-1}^{crop},\mathcal{P}_{t}^{crop},\mathcal{B}_{t-1}))),$ (1)
@@ 3.4 Proposed P2P Framework
$[\Delta x_{t},\Delta y_{t},\Delta z_{t},\Delta\theta_{t}]=F_{m}^{pp}(\mathcal{P}_{t-1}^{crop},\mathcal{P}_{t}^{crop},\mathcal{B}_{t-1}),$ (2)
where $F_{m}^{pp}$ can mathematically express our P2P framework. Its overall architecture is illustrated in Fig. 4. The framework consists of three components: feature extractor, part-to-part motion modeling, and motion prediction head. The feature extractor distinguishes and extracts spatial structure features of $\mathcal{P}_{t-1}^{crop}$ and $\mathcal{P}_{t}^{crop}$ by a weight-shared network using dual branches. The part-to-part motion modeling module first fuses each corresponding part between the spatial structures of $\mathcal{P}_{t-1}^{crop}$ and $\mathcal{P}_{t}^{crop}$, and then models target-related motion cues in each part. Finally, the prediction head infers the 4-DOF relative motion.
@@ 3.6 Version II: P2P-voxel
$\begin{array}[]{c}1\times 4\end{array}$ $\begin{bmatrix}\!1,512,1\!\\ \!1,256,1\!\\ \!1,128,1\!\\ \!1,4,1\!\\ \end{bmatrix}\times 1$ $\begin{array}[]{c}1\times 4\end{array}$ $\begin{bmatrix}\!1,512,1\!\\ \!1,256,1\!\\ \!1,128,1\!\\ \!1,4,1\!\\ \end{bmatrix}\times 1$
@@ 3.7 Effectiveness Analysis
Deep Tracking Framework Recent years, the appearance matching [11, 31] framework has attracted significant attention. However, its template matching mechanism tend to interfere with target’s motion cues, thereby requiring complex structures for precise target localization. For example, the most recent appearance matching based trackers CXTrack [48] and MBPTrack [49] achieve advanced performance contributed to a series of prior structural designs, including a target-centric transformer, a box-prior localization network and so on. The motion-centric framework, represented by M2Track [54], appears to be a more promising approach. It fully exploits the motion cues of the target between consecutive frames, achieving state-of-the-art performance with an interpretable pipeline. Even better, our proposed P2P framework integrates part-to-part motion modeling, enabling the direct inference of accurate relative motion without relying on architectural components, such as segmentation module, two-stage design found in M2Track. The effectiveness and potential of fusing each corresponding part between spatial structures of consecutive point clouds have been demonstrated in the experiment section. More details are as follows.
@@
For P2P-point, $n$ is $C$, and each part feature $\in 1\times 1$, where part-to-part refers to channel-to-channel. For P2P-voxel, $n$ is $W\cdot L$, and each part feature $\in 1\times C$, where part-to-part refers to spatial-to-spatial. As a result, more fine-grained motion cues in each part are modeled, guiding a more accurate and robust motion prediction. Benefiting from this, our methods work in a one-stage manner without extra segmentation and two-stage refinement modules used in M2Track [54], while demonstrating state-of-the-art performance and high speed.
P2P-point v.s. P2P-voxel. The superior performance of the proposed framework stems from its part-to-part motion modeling, which facilitates robust and accurate motion cues for tracking, as demonstrated in our manuscript. A critical component of this process is part-to-part fusion, which directly impacts the quality of motion cues and thus affects the overall performance. Compared to P2P-point, P2P-voxel has inherent advantages in fusing corresponding parts between consecutive two-frame point clouds. Through voxel-based representation, P2P-voxel preserves the spatial structure of point clouds, enabling explicit part-to-part fusion, where each pixel explicitly represents a distinct part. This precise correspondence allows for more accurate motion modeling. In contrast, P2P-point utilizes a point-based representation, relying on pooling operations to maintain the permutation invariance of point clouds. This approach embeds the spatial structure into a 1D semantic feature, where each channel implicitly represents a part of the overall point cloud information. As a result, P2P-point achieves implicit part-to-part fusion between consecutive point clouds. While both point and voxel representation-based models have their merits, explicit part-to-part fusion used in P2P-voxel allows for a more precise alignment and interaction between corresponding parts of consecutive point clouds, thereby outperforming the implicit approach of P2P-point. Therefore, P2P-voxel presents performance advantage over P2P-point.
@@
Table 2: Backbone architectures of our customized PointNet and VoxelNet. “$[k,c,s]\times n$” means $n$ 1D convolution layers with kernel size $k$, output channel $c$ and stride $s$, while “$[k\times k\times k,c,s]\times n$” means $n$ 3D sparse convolution layers with kernel size $k\times k\times k$, output channel $c$ and stride $s$. “#$n$” represents the $n$-th stage within the backbones.
Output Size Point. Embed Output Size Voxel. Embed
#1 $\begin{array}[]{c}1024\times 64\end{array}$ $\begin{bmatrix}\!1,64,1\!\\ \end{bmatrix}\times 2$ $\begin{array}[]{c}64\times 64\times 10\times 32\end{array}$
$\begin{bmatrix}\!3\times 3\times 3,16,1\!\\ \end{bmatrix}\times 2$
$\begin{bmatrix}\!3\times 3\times 3,32,1\!\\ \end{bmatrix}\times 1$
#2 $\begin{array}[]{c}1024\times 128\end{array}$ $\begin{bmatrix}\!1,128,1\!\\ \end{bmatrix}\times 1$ $\begin{array}[]{c}32\times 32\times 5\times 64\end{array}$
$\begin{bmatrix}\!3\times 3\times 3,32,1\!\\ \end{bmatrix}\times 2$
$\begin{bmatrix}\!3\times 3\times 3,64,1\!\\ \end{bmatrix}\times 1$
#3 $\begin{array}[]{c}1024\times 1024\end{array}$ $\begin{bmatrix}\!1,1024,1\!\\ \end{bmatrix}\times 1$ $\begin{array}[]{c}16\times 16\times 3\times 128\end{array}$
$\begin{bmatrix}\!3\times 3\times 3,64,1\!\\ \end{bmatrix}\times 2$
$\begin{bmatrix}\!3\times 3\times 3,128,1\!\\ \end{bmatrix}\times 1$
$\begin{array}[]{c}1\times 1024\end{array}$ Maxpooling $\begin{array}[]{c}16\times 16\times 128\end{array}$ Flatten & Linear
@@ 4 Experiment
4.1 Implementation Details
Model Hyper-parameters. We define search regions for two consecutive frames at timestamps $t-1$ and $t$ by centering on the previous target bounding box within a 3D space. The ranges are defined as [(-4.8,4.8),(-4.8,4.8),(-1.5,1.5)] and [(-1.92,1.92),(-1.92,1.92),(-1.5,1.5)] to contain relevant points for cars and humans, respectively. To be a fair comparison, we follow existing methods [31, 54] to sub-sample 1024 points using farthest point sampling [29] within the derived search regions as inputs for P2P-point. For P2P-voxel, we do not down-sample point clouds. Instead, we divide the search regions into 128$\times$128$\times$20 voxel grids, which serve as inputs for the model. More model details can be found in Tab. 1.
Tab. 2 presents the backbone architectures utilized for P2P-point and P2P-voxel, i.e., “Point. Embed” and “Voxel. Embed” in Tab. 1of the main manuscript. We modify the PointNet [28] and VoxelNet [59] as our backbones to extract features. Specifically, we remove task layers and T-Net of PointNet, using only 1D convolution layers with batch normalization (BN) and rectified linear unit (ReLU) function to extract point features. Similarly, we remove the task layers of VoxelNet and adopt 3D sparse convolution blocks without residual structure to extract the features of input voxels.
@@ 4.1 Implementation Details
Input: A $T$-frame point cloud sequence ${\{\mathcal{P}_{t}\}_{t=1}^{T}}$; A target bounding box $\mathcal{B}_{1}=(x_{1},y_{1},z_{1},w_{1},h_{1},l_{1},\theta_{1})$ in the $1$-st frame of the given sequence.
Output: A series of predicted bounding boxes $\{\mathcal{B}_{t}=(x_{t},y_{t},z_{t},w_{t},h_{t},l_{t},\theta_{t})\}_{t=2}^{T}$ from $2$-nd frame to $T$-th frame.
1 for $t=2$ to $T$ do
2       Crop previous search region $\mathcal{P}_{t-1}^{crop}\in\mathbb{R}^{1024\times 3}$ from point cloud $\mathcal{P}_{t-1}$ in the ($t-1$)-th frame;
3       Crop current search region $\mathcal{P}_{t}^{crop}\in\mathbb{R}^{1024\times 3}$ from point cloud $\mathcal{P}_{t}$ in the $t$-th frame;
4       Use a weight-shared PointNet to encode $\mathcal{P}_{t-1}^{crop}$ and $\mathcal{P}_{t}^{crop}$ to embeddings $\mathcal{F}_{t-1}\in\mathbb{R}^{1\times 1024}$ and $\mathcal{F}_{t}\in\mathbb{R}^{1\times 1024}$;
5       Fuse $\mathcal{F}_{t-1}$ and $\mathcal{F}_{t}$ in a spatial-aligned manner to obtain fused feature $\mathcal{F}_{fusion}^{sa}\in\mathbb{R}^{256\times 1024}$;
6       Obtain 4-DOF relative motion $\mathcal{M}_{t-1,t}=(\Delta x_{t},\Delta y_{t},\Delta z_{t},\Delta\theta_{t})$ by applying a MLP based prediction head;
7       Transform previous $\mathcal{B}_{t-1}$ using $\mathcal{M}_{t-1,t}$ to output $\mathcal{B}_{t}$ in current frame.
8
Return $\{\mathcal{B}_{t}\}_{t=2}^{T}$.
Algorithm 1 Tracking process of P2P-point
Input: A $T$-frame point cloud sequence ${\{\mathcal{P}_{t}\}_{t=1}^{T}}$; A target bounding box $\mathcal{B}_{1}=(x_{1},y_{1},z_{1},w_{1},h_{1},l_{1},\theta_{1})$ in the $1$-st frame of the given sequence.
Output: A series of predicted bounding boxes $\{\mathcal{B}_{t}=(x_{t},y_{t},z_{t},w_{t},h_{t},l_{t},\theta_{t})\}_{t=2}^{T}$ from $2$-nd frame to $T$-th frame.
1 for $t=2$ to $T$ do
2       Crop previous search region $\mathcal{V}_{t-1}^{crop}\in\mathbb{R}^{128\times 128\times 20}$ from point cloud $\mathcal{P}_{t-1}$ in the ($t-1$)-th frame;
3       Crop current search region $\mathcal{V}_{t}^{crop}\in\mathbb{R}^{128\times 128\times 20}$ from point cloud $\mathcal{P}_{t}$ in the $t$-th frame;
4       Use a weight-shared VoxelNet to encode $\mathcal{V}_{t-1}^{crop}$ and $\mathcal{V}_{t}^{crop}$ to $\mathcal{F}_{t-1}\in\mathbb{R}^{16\times 16\times 128}$ and $\mathcal{F}_{t}\in\mathbb{R}^{16\times 16\times 128}$;
5       Fuse $\mathcal{F}_{t-1}$ and $\mathcal{F}_{t}$ in a spatial-aligned manner to obtain fused feature $\mathcal{F}_{fusion}^{sa}\in\mathbb{R}^{4\times 4\times 1024}$;
6       Obtain 4-DOF relative motion $\mathcal{M}_{t-1,t}=(\Delta x_{t},\Delta y_{t},\Delta z_{t},\Delta\theta_{t})$ by applying a 2D Conv based prediction head;
7       Transform previous $\mathcal{B}_{t-1}$ using $\mathcal{M}_{t-1,t}$ to output $\mathcal{B}_{t}$ in current frame.
8
Return $\{\mathcal{B}_{t}\}_{t=2}^{T}$.
Algorithm 2 Tracking process of P2P-voxel
@@
4.2 Experiment Setting
Datasets. We conduct comprehensive experiments on three widely used datasets, including KITTI [10], NuScenes [2] and Waymo Open Dataset (WOD) [34]. KITTI consists of 21 training and 29 test point cloud sequences, respectively. Due to the unavailability of test labels, we divide the training sequences into training set [0-17), validation set [17-19) and test set [19-21). Compared to KITTI, NuScenes and WOD provide more challenging and large-scale scenes. NuScenes contains 700 and 150 scenes for training and testing, while WOD comprises 1121 tracklets categorized into easy, medium and hard sub-sets based on the point cloud sparsity.